# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ I.2 Ion Coulomb crystals in a laboratory
Confinement of ions in higher order rf traps leads to new possibilities and challenges. A good review on trapping of charged particles in multipole rf traps is given in Ref.[120]. Rf traps of order 2qq leads in general to effective pseudo potential of order 2qq-2, and hence generally a much shallower potential than an harmonic potential, and less micromotion at positions close to the trap center. This latter feature has for decades been exploited in experiments involving buffer gas cooling of dilute ensembles of molecular ions [121]. In the context of Coulomb crystallization multipole traps have the prospect of leading to three dimensional crystals with non-uniform ion densities, something that more recently have been investigated theoretically [52] and pursued experimentally [238]. The current lack of high quality experimental results relies mainly on the ability to operate perfect multipole traps. Any lower order perturbations will significantly distort the effective potential with the result of more local effective potential minima [245]. More fundamentally, it is still unclear what effect the anharmonicity in the effective potential will have on rf induced heating of dense ion ensembles.
Figure 6: Cross-sectional sketch of Penning trap that employs cylindrical electrodes, similar to that used in [37]. The blue arrows represent Doppler laser cooling beams that cool ion motion parallel and perpendicular to the magnetic field. The blue spheroid is an ion crystal in the center of the trap. Not shown are light collection optics that image ion fluorescence both parallel and perpendicular to the magnetic field. The central electrode is segmented to enable the application of a rotating wall potential. See the text.
@@ I.2.2 Penning trap
In contrast to rf traps, Penning traps (and the closely related Penning-Malmberg traps) employ both static electric and magnetic fields to confine charged particles [98, 326]. A set of electrodes, frequently a co-axially aligned array of cylindrical electrodes as sketched in Fig. 6, generates an electrostatic trap potential $\phi_{T}$ that has symmetry about the common axis of the cylinders, denoted the $\hat{z}-$axis. With a positive potential applied to the end cylinders relative to the central cylinder, $\phi_{T}$ provides axial confinement for positively charged ions. However, ${\nabla}^{2}\phi_{T}=0$ implies the electric field near the center of the Penning trap is anti-confining in the radial direction. Radial confinement is then obtained by immersing the trap in a strong uniform magnetic field $^{\mathbf{B}}=B\hat{z}$ directed along the symmetry axis of the trap.
In more detail, ions in the trap undergo an ${\mathbf{E}}\times{\mathbf{B}}$ drift, which, due to the radial electric field of the trapping potential, is a rotation of the ion crystal about the symmetry axis of the trap. Rotation in the presence of the strong axial magnetic field produces a Lorentz force directed radially inward, providing radial confinement. Theoretical and experimental work shows that for a trap with axial symmetry and in the absence of external forces, a collection of charged particles confined in a Penning trap evolves to a thermal equilibrium state characterized by rigid rotation [98, 68, 89, 34]. Throughout this manuscript we use $\omega_{r}$ to denote the rotation frequency of the ion crystal. Note that the energy of an ensemble of ions in a Penning trap is lowered when the ions in the crystal move radially outward. Radial confinement in a Penning trap is therefore not an energetic confinement. It can be understood in terms of the conservation of another conserved quantity, the canonical angular momentum about the symmetry axis of the trap [98]. This motivates the construction of Penning traps with good axial symmetry.
@@
where $z$ and $\rho$ are axial and cylindrical coordinates, $m$ is the mass and $Q$ is the charge of the ion, and the trap frequency $\omega_{z}$ characterizes the strength of the axial confinement. It depends on the potential difference $V_{T}$ applied between the end and central electrodes as
$\omega_{z}=\sqrt{\frac{QV_{0}}{md_{T}^{2}}}\>.$ (21)
Here $V_{0}$ is the voltage applied between the end and central cylinders and $d_{T}$ is a characteristic distance describing the trap. The volume over which Eq. (20) is a good approximation can be increased by employing trap electrodes with hyperboloid shapes [39].
It is frequently convenient to describe the dynamics of an ion crystal in a Penning trap in a frame rotating with the ion crystal rotation frequency $\omega_{r}$. In this rotating frame, the radially deconfining trap potential of Eq. (20) turns into an effective, radially confining potential
@@
where $\beta=\frac{\omega_{r}\left(\Omega_{c}-\omega_{r}\right)}{\omega_{z}^{2}}-\frac{1}{2}$. Here $\Omega_{c}=\frac{QB}{m}$ is the ion cyclotron frequency and the term proportional to $\omega_{r}\Omega_{c}$ is from the Lorentz force generated by the rotation of the ion crystal in the presence of the strong magnetic field. The parameter $\beta$, sometimes called the trap anisotropy parameter, characterizes the relative strength of the radial confinement compared to the axial confinement.
So-called Penning-Malmberg traps [74, 75] are employed in non-neutral plasma physics for studying non-equilibrium dynamics and the transport of particles and energy across magnetic field lines, and for the long-term confinement of anti-matter [230, 6, 113]. Here the central cylindrical section of the trap (the segmented central electrode in Fig. 6) is a cylinder whose length is large compared to its diameter. In Penning-Malmberg traps the number of trapped charged particles can be large ($10^{8}$), filling a volume where Eq. (20) is no longer a good approximation. For a discussion of other configurations of trap electrodes, see [326].
@@
where $\epsilon_{0}$ is the permittivity of the vacuum. In contrast to rf traps, this result is independent of the details of the trapping potential $\phi_{T}$. Specifically, even if the trapping potential (Eq. (20)) contains higher order terms ($z^{4}$, $\rho^{4}$, $z^{2}\rho^{2}$, $\ldots$), the 3D ion density of the crystal is constant and given by Eq. (23). The density is maximized at $\omega_{r}=\Omega_{c}/2$, known as the Brillouin density limit [36]. Typical densities $n_{0}$ demonstrated in the laboratory range between $10^{13}$ m-3 to $10^{16}$ m-3.
@@
where $\omega_{m}=\frac{\Omega_{c}}{2}-\sqrt{(\frac{\Omega_{c}}{2})^{2}-\frac{\omega_{z}^{2}}{2}}$ is usually called the magnetron frequency. It is the frequency at which a single trapped ion undergoes a circular motion about the center of the trap due to ${\mathbf{E}}\times{\mathbf{B}}$ drift. In addition, in the limit that the uniform density ion crystal can be treated as a uniformly charged fluid (valid for ion crystals with dimensions more than a few interparticle spacings), the low temperature equilibrium shape of the fluid (or crystal) is a uniform density spheroid with an aspect ratio $\alpha=Z_{p}/R_{p}$ determined by $\omega_{r}$, $\Omega_{c}$ and $\omega_{z}$ through the parameter $\beta$. Here $2Z_{p}$ is the axial extent and $2R_{p}$ the diameter of the spheroid. The detailed functional dependence is discussed in Refs. [28, 98] and is in good agreement with experimental measurements [34, 28].
Figure 7: Simple qualitative sketch of the ion number density $n_{0}$ and different spheroidal crystal shapes obtained as a function of the ion crystal rotation frequency $\omega_{r}$ in a Penning ion trap. The maximum density $n_{B}$, called the Brillouin limit, occurs for $\omega_{r}=\Omega_{c}/2$.
Figure 7 illustrates the qualitative dependence of the ion crystal density and shapes obtained as a function of the rotation frequency $\omega_{r}$ for fixed $\omega_{z}$ and $\Omega_{c}$. For $\omega_{r}$ slightly greater than $\omega_{m}$, the crystal consists of a single plane corresponding to $\alpha\approx 0$. As $\omega_{r}$ is increased, the radius of the ion crystal decreases due to the increasing radial confinement and at some rotation frequency ions move out of the $z=0$ plane, forming a 3-dimensional crystal. Further increases in $\omega_{r}$ increases the aspect ratio $\alpha$ as well as the ion density $n_{0}$. The largest $\alpha$ and highest density $n_{0}$, known as the Brillouin limit, is obtained for $\omega_{r}=\Omega_{c}/2$. For $\omega_{r}>\Omega_{c}/2$, the centrifugal acceleration weakens the radial confinement, resulting in a decreasing ion crystal aspect ratio and ion density. For ion crystals with small numbers of ions, the radial confinement with $\omega_{r}\sim\Omega_{c}/2$ can be made sufficiently strong compared to the axial confinement that the ions form a 1-dimensional crystal along the symmetry axis of the trap [203]. These structural phase transitions ($2D\rightarrow 3D$ [214] and $3D\rightarrow 1D$ [203]) have been observed in the lab.
@@
Here the frequency $\omega_{W}$ characterizes the strength of the quadrupole potential and $x_{\text{rot}}$ and $y_{\text{rot}}$ are rotating frame coordinates along the strong and weak axes of the rotating quadrupole potential. The rotating quadrupole potential breaks the symmetry of the trap potential and changes the shape of the ion crystal, providing a restoring force that controls the azimuthal orientation of the ion crystal in the frame of the rotating potential.
@@ I.3 Laser cooling
The low temperatures required for the formation of the ion Coulomb crystal is typically reached by means of laser cooling [58, 61, 248]. This mechanism is based on scattering of laser light and uses the mechanical effects of radiation to transfer energy from the ions’ motion to the modes of the electromagnetic field. As such, the stationary state of the ions results from the dynamical equilibrium of scattering processes that heat and that cool the motion. The width of the stationary energy distribution is determined by the characteristic linewidth of the scattering processes and is (often improperly) associated with a temperature. Laser cooling allows to prepare ions at temperatures ranging from hundreds of millikelvin down to fractions of a microkelvin, depending on the technique which is employed and on the ion species.
@@
Laser cooling of ion crystals employs, typically, techniques developed for single ions. In fact, multiple photon scattering events are very rare because the ion density is sufficiently low that the ion crystals are optically thin. Hence each ion is effectively an independent scatterer [222, 221]. Nevertheless, there are some challenges to consider when applying laser cooling protocols to ion ensembles. One important challenge is that processes that heat in general become increasingly important with the number of ions in the crystal. Another major challenge consists in the fact that the non-inertial forces of the trap cannot be neglected for the ions located away from the symmetry axes. This makes it difficult to warrant that the sample is uniformly cooled at the same temperature. We will discuss the implications on studies of thermalization in Sec. III. In what follows we provide a brief review of the state-of-the art in preparing ion ensembles at low temperatures. For a general review of laser cooling with trapped ions, we refer the reader to [106].
@@
The most commonly used laser cooling technique is Doppler laser cooling, which can produce temperatures on the order of 1 mK or lower [106, 302, 333]. Doppler laser cooling has been used to prepare one-dimensional (1D), linear chain crystals of up to $\sim 100$ ions in linear rf traps (see Sec. II.1) and two-dimensional (2D), single plane crystals of ions of up to $\sim 500$ ions in both Penning and rf ion traps (see Sec. II.2). Much larger three-dimensional (3D) ion crystals ($N>10^{5}$) have been formed through Doppler laser cooling in both Penning and rf ion traps (see Sec. II.3).
@@
Once formed, 1D and 2D ion crystals have been cooled to temperatures well below the mK regime by means of so-called sub-Doppler cooling techniques. Examples of sub-Doppler cooling techniques include polarization-gradient cooling [23, 164, 193], sideband cooling [77, 219] and Electromagnetically Induced Transparency (EIT) cooling [223, 265]. For completeness we also mention protocols for simultaneously cooling the vibrational modes of an ion chain to the ground state that are based on designing the chain spectrum by means of an external field [112, 336]. These sub-Doppler cooling techniques can prepare the crystal close to the zero-point motion. Sideband cooling is routinely used to sequentially cool the vibrational modes of an ion chain in a linear Paul trap to near the quantum mechanical ground state state [106]. Moreover, in Penning traps, small crystals with less than 10 ions have been successfully sideband-cooled [304].
@@
EIT cooling has been demonstrated to be particularly efficient for cooling crystals of ions to the ground state, thanks to the Fano-like profile of the scattering processes [220], which permits to cool simultaneously several vibrational modes of the crystal [284]. In a linear Paul trap, ion chains ranging from a few ions [196] to the transverse motion of a chain of 18 ions [189] and 40 ions [108] have been EIT cooled to near the ground state. It has also been employed to simultaneously cool the out-of-plane modes with single plane crystals of a few hundred ions to near the ground state of the vibrational motion in Penning traps [162, 290] and rf traps [171, 129].
@@
While 1D and 2D crystal formation has been limited to a few hundred ions, Doppler laser cooling has been used to prepare much larger, mesoscopic, 3D crystals of $N>10^{5}$ ions in both Penning and rf traps. Laser cooling may play a role in the size of ion crystals that can be produced in the lab. Below we discuss experimental demonstrations of some of the largest trapped ion crystals that have been prepared in the lab.
@@
Penning traps employ static confining fields and the non-thermal motion of ions in the trap is a rigid body rotation that is not readily shared with the thermal degrees of freedom. Therefore, Doppler cooling of rotating ion ensembles in Penning traps has been relatively straightforward, although Doppler shifts due the ion crystal rotation can produce temperatures that are slightly elevated from standard Doppler cooling limits [318]. Three-dimensional crystals with $N\sim 5\times 10^{5}$ ions have been Doppler cooled in Penning traps to milliKelvin temperatures ($\Gamma>200$), where a body centered cubic (bcc) lattice structure in the crystal’s interior was observed [155, 308]. Doppler cooling has been demonstrated to be effective for larger ensembles: Nonneutral plasmas of $N\sim 10^{8}$ Mg+ ions in Penning-Malmberg traps have been prepared at temperatures less than 50 mK [3, 4]. Here the ions are strongly correlated ($\Gamma\sim 20$) but above the temperature for crystal formation. The lack of crystal formation is not understood, but processes that heat the ions and increase the ion temperature in general increase with $N$. For example, the potential energy stored in the ion crystal increases with $N$ and in Penning traps small imperfections in the construction of the trap can drive a slow radial expansion of the ion crystal that converts potential energy to kinetic energy [159, 98].
@@
Cooling large ion ensembles in rf traps is challenging due to the time-dependent potential. In fact, the rf-driven micromoton can be converted into thermal motion by the Coulomb interactions, heating the crystal. This heating rate increases as the temperature is lowered and the number of ions increases, giving rise to chaotic dynamics [26, 232] and limiting the cooling efficiency. Therefore, the formation of large ion crystals in rf traps requires large laser cooling powers. Despite these difficulties, large crystals composed of $10^{4}$ to $10^{5}$ ions have been prepared in rf traps [84, 146]. It has been observed that once the ion crystal forms, the rf heating can become significantly reduced. A lower limit on the Coulomb coupling parameter of $\Gamma\gtrsim 70$ has been estimated from the observed shell structure of these large ion crystals in rf traps. Direct measurements of the ion temperature are challenging with 3D crystals in rf traps because of the large driven micromotion.

### v2

@@ I.2 Ion Coulomb crystals in a laboratory
Confinement of ions in higher order rf traps leads to new possibilities and challenges. A good review on trapping of charged particles in multipole rf traps is given in Ref. [130]. Rf traps of order 2ll leads in general to effective pseudo potential of order 2ll-2, and hence to a much shallower potential than the harmonic potential, as well as to less micromotion at positions close to the trap center. This latter feature has been exploited for decades in experiments involving buffer gas cooling of dilute ensembles of molecular ions [131]. In the context of Coulomb crystallization multipole traps are expected to realize three-dimensional crystals with non-uniform ion densities, see the theoretical studies of [56, 57] and the experimental realization of [265]. The current lack of high quality experimental results is mainly due to the limitations in the ability to operate perfect multipole traps. Lower order perturbations significantly distort the effective potential with the result of more local effective potential minima [272]. It is still unclear what is the effect of the anharmonicity in the effective potential on rf induced heating for dense ion ensembles.
@@ I.2.2 Penning trap
In contrast to rf traps, Penning traps (and the closely related Penning-Malmberg traps) employ both static electric and magnetic fields to confine charged particles [105, 362]. A set of electrodes, frequently a co-axially aligned array of cylindrical electrodes as sketched in Fig. 6, generates an electrostatic trap potential $\phi_{T}$ that has symmetry about the common axis of the cylinders, denoted by the $\hat{z}-$axis. With a positive potential applied to the end cylinders relative to the central cylinder, $\phi_{T}$ provides axial confinement for positively charged ions. However, the Laplace equation ${\nabla}^{2}\phi_{T}=0$ implies the electric field near the center of the Penning trap is anti-confining in the radial direction. Radial confinement is then obtained by immersing the trap in a strong uniform magnetic field $^{\mathbf{B}}=B\hat{z}$ directed along the symmetry axis of the trap.
Figure 6: Cross-sectional sketch of a Penning trap that employs cylindrical electrodes, similar to that used in [39]. The thick arrows represent Doppler laser cooling beams that cool ion motion parallel and perpendicular to the magnetic field. The spheroid they are pointing at is an ion crystal in the center of the trap. For trapping of positively charged ions, the outer electrodes are biased with a positive potential with respect to the central electrode, typically with $V_{2}>V_{1}>0$. The central electrode is segmented to enable the application of a rotating wall potential. See the text. Not shown are light collection optics that image ion fluorescence both parallel and perpendicular to the magnetic field.
In more detail, ions in the trap undergo an ${\mathbf{E}}\times{\mathbf{B}}$ drift, which, due to the radial electric field of the trapping potential, is a rotation of the ion crystal about the symmetry axis of the trap. Rotation in the presence of the strong axial magnetic field produces a Lorentz force directed radially inward, providing radial confinement. Theoretical and experimental work shows that for a trap with axial symmetry and in the absence of external forces, a collection of charged particles confined in a Penning trap evolves to a thermal equilibrium state characterized by rigid rotation [37, 95, 105, 73]. Throughout this manuscript we use $\omega_{r}$ to denote the rotation frequency of the ion crystal. Note that the energy of an ensemble of ions in a Penning trap is lowered when the ions in the crystal move radially outward. Radial confinement in a Penning trap is therefore not an energetic confinement. It can be understood in terms of the conservation of the canonical angular momentum about the symmetry axis of the trap [105]. This motivates the construction of Penning traps with good axial symmetry.
@@
where $z$ and $\rho$ are axial and cylindrical coordinates, $m$ is the mass and $Q$ is the charge of the ion. The trap frequency $\omega_{z}$ characterizes the strength of the axial confinement: it scales as the square root of the potential difference between the end and central electrodes and inversely with the size of the trap. For a trap with two end electrodes and one central electrode,
$\omega_{z}=\sqrt{\frac{QV_{0}}{md_{T}^{2}}}\>,$ (21)
where $V_{0}$ is the voltage applied between the end and central cylinders and $d_{T}$ is a characteristic distance describing the trap. The volume over which Eq. (20) is a good approximation can be increased by employing trap electrodes with hyperboloid shapes [41].
It is frequently convenient to describe the dynamics of an ion crystal in a Penning trap in a frame rotating with rotation frequency $\omega_{r}$ of the ion crystal. In this rotating frame, the radially deconfining trap potential of Eq. (20) turns into an effective, radially confining potential
@@
where $\beta=\frac{\omega_{r}\left(\Omega_{c}-\omega_{r}\right)}{\omega_{z}^{2}}-\frac{1}{2}$. Here, $\Omega_{c}=\frac{QB}{m}$ is the ion cyclotron frequency and the term proportional to $\omega_{r}\Omega_{c}$ is due to the Lorentz force generated by the rotation of the ion crystal in the presence of the strong magnetic field. The parameter $\beta$, sometimes called the trap anisotropy parameter, characterizes the relative strength of the radial confinement compared to the axial confinement.
So-called Penning-Malmberg traps [79, 80] are employed in non-neutral plasma physics for studying non-equilibrium dynamics and the transport of particles and energy across magnetic field lines, and for the long-term confinement of anti-matter [256, 6, 122]. Here, the central cylindrical section of the trap (the segmented central electrode in Fig. 6) is a cylinder whose length is large compared to its diameter. In Penning-Malmberg traps the number of trapped charged particles can be large ($10^{8}$), filling a volume where Eq. (20) is no longer a good approximation. For a discussion of other configurations of trap electrodes, see [362].
@@
where $\epsilon_{0}$ is the permittivity of the vacuum. In contrast to rf traps, this result is independent of the details of the trapping potential $\phi_{T}$. Specifically, even if the trapping potential (Eq. (20)) contains higher order terms ($z^{4}$, $\rho^{4}$, $z^{2}\rho^{2}$, $\ldots$), the three-dimensional density of the ion crystal is constant and given by Eq. (23). The density is maximized at $\omega_{r}=\Omega_{c}/2$, known as the Brillouin density limit [38]. Typical densities $n_{0}$ demonstrated in the laboratory range between $10^{13}$ m-3 to $10^{16}$ m-3 [163, 105].
@@
where $\omega_{m}=\frac{\Omega_{c}}{2}-\sqrt{(\frac{\Omega_{c}}{2})^{2}-\frac{\omega_{z}^{2}}{2}}$ is usually called the magnetron frequency. It is the frequency at which a single trapped ion undergoes a circular motion about the center of the trap due to ${\mathbf{E}}\times{\mathbf{B}}$ drift. In addition, in the limit that the uniform density ion crystal can be treated as a uniformly charged fluid (valid for ion crystals with dimensions more than a few interparticle spacings), the low temperature equilibrium shape of the fluid (or crystal) is a uniform density spheroid with an aspect ratio $\alpha=Z_{p}/R_{p}$ determined by $\omega_{r}$, $\Omega_{c}$ and $\omega_{z}$ through the parameter $\beta$. Here, $2Z_{p}$ is the axial extent and $2R_{p}$ the diameter of the spheroid. The detailed functional dependence is discussed in Refs. [31, 105] and is in good agreement with experimental measurements [37, 31].
Figure 7: Simple qualitative sketch of the ion number density $n_{0}$ and different spheroidal crystal shapes obtained as a function of the ion crystal rotation frequency $\omega_{r}$ in a Penning trap. The maximum density $n_{B}$, called the Brillouin limit, occurs for $\omega_{r}=\Omega_{c}/2$.
Figure 7 illustrates the qualitative dependence of the ion crystal density and shapes obtained as a function of the rotation frequency $\omega_{r}$ for fixed $\omega_{z}$ and $\Omega_{c}$. For $\omega_{r}$ slightly greater than $\omega_{m}$, the crystal consists of a single plane corresponding to $\alpha\approx 0$. As $\omega_{r}$ is increased, the radius of the ion crystal decreases due to the increasing radial confinement and at some rotation frequency ions move out of the $z=0$ plane, forming a 3-dimensional crystal. Further increases in $\omega_{r}$ increases the aspect ratio $\alpha$ as well as the ion density $n_{0}$. The largest $\alpha$ and highest density $n_{0}$, known as the Brillouin limit, is obtained for $\omega_{r}=\Omega_{c}/2$. For $\omega_{r}>\Omega_{c}/2$, the centrifugal acceleration weakens the radial confinement, resulting in a decreasing ion crystal aspect ratio and ion density. For ion crystals with small numbers of ions, the radial confinement with $\omega_{r}\sim\Omega_{c}/2$ can be made sufficiently strong compared to the axial confinement that the ions form a 1-dimensional crystal along the symmetry axis of the trap [226]. These structural phase transitions (from two dimensions to three dimensions [238] and from three dimensions to one dimension [226]) have been observed in the lab.
@@
Here, the frequency $\omega_{W}$ characterizes the strength of the quadrupole potential and $x_{\text{rot}}$ and $y_{\text{rot}}$ are rotating frame coordinates along the strong and weak axes of the rotating quadrupole potential. The rotating quadrupole potential breaks the symmetry of the trap potential and changes the shape of the ion crystal, providing a restoring force that controls the azimuthal orientation of the ion crystal in the frame of the rotating potential.
@@ I.3 Laser cooling
The low temperature required for the formation of the ion Coulomb crystal is typically reached by means of laser cooling [63, 66, 275]. This mechanism is based on scattering of laser light and uses the mechanical effects of radiation to transfer energy from the ions’ motion to the modes of the electromagnetic field. As such, the stationary state of the ions results from the dynamical equilibrium of scattering processes that heat and that cool the motion. The width of the stationary energy distribution is determined by the characteristic linewidth of the scattering processes and is often associated with a temperature. See Sec. III.1 for a detailed discussion. Laser cooling allows to prepare ions at temperatures ranging from hundreds of millikelvin down to fractions of a microkelvin, depending on the technique which is employed and on the ion species.
@@
Laser cooling of ion crystals employs, typically, techniques appropriate for cooling individual ions. In fact, multiple photon scattering events are very rare because the ion crystals are optically thin. In this limit, each ion of the crystal is effectively an independent scatterer [246, 247]. Nevertheless, there are some challenges to consider when applying laser cooling protocols to ion ensembles. One important challenge is that processes that heat in general become increasingly important with the number of ions in the crystal. Another major challenge consists in the fact that the non-inertial forces of the trap cannot be neglected for the ions located away from the symmetry axes. This makes it difficult to ensure that the sample is uniformly cooled at the same temperature. We will discuss the implications on studies of thermalization in Sec. III. In what follows we provide a brief review of the state-of-the art in preparing ion ensembles at low temperatures. For a comprehensive review of laser cooling with trapped ions, we refer the reader to [113].
@@
The most commonly used laser cooling technique is Doppler laser cooling, which can produce temperatures on the order of 1 mK or lower [369, 337, 113]. Doppler laser cooling has been used to prepare one-dimensional (1D), linear chain crystals of up to $\sim 100$ ions in linear rf traps (see Sec. II.1) and two-dimensional (2D), single-plane crystals of ions of up to $\sim 500$ ions in both Penning and rf ion traps (see Sec. II.2). Much larger three-dimensional (3D) ion crystals have been formed through Doppler laser cooling in both Penning ($10^{5}$ ions) and rf ion traps ($10^{4}$ ions) (see Sec. II.3).
@@
Once formed, 1D and 2D ion crystals are cooled to temperatures well below the mK regime by means of so-called sub-Doppler cooling techniques. Examples of sub-Doppler cooling techniques include polarization-gradient cooling [26, 177, 212], sideband cooling [83, 244] and Electromagnetically Induced Transparency (EIT) cooling [248, 297]. For completeness we also mention protocols for simultaneously cooling the vibrational modes of an ion chain to the ground state that are based on designing the chain spectrum by means of an external field [373, 121]. These sub-Doppler cooling techniques can prepare the crystal close to the zero-point motion. Sideband cooling is routinely used to sequentially cool the vibrational modes of an ion chain in a linear Paul trap to near the quantum mechanical ground state state [113]. Moreover, in Penning traps, small crystals with less than 10 ions have been successfully sideband-cooled [339].
@@
EIT cooling has been demonstrated to be particularly efficient for cooling crystals of ions to the ground state, thanks to the Fano-like profile of the scattering processes [245], which permits to cool simultaneously several vibrational modes of the crystal [318]. In a linear Paul trap, ion chains ranging from a few ions [215] to the transverse motion of a chain of 18 ions [206] and 40 ions [117] have been EIT cooled to near the ground state. EIT cooling has also been employed to simultaneously cool the out-of-plane modes of single-plane crystals to near the ground state of the vibrational motion in Penning traps  [175, 325] and rf traps [186, 140].
@@
While 1D and 2D crystal formation has been limited to a few hundred ions, Doppler laser cooling has been used to prepare much larger, mesoscopic, 3D crystals in both Penning and rf traps. Laser cooling may play a role in determining the size of ion crystals that can be produced in the lab. Below we discuss experimental demonstrations of some of the largest trapped ion crystals that have been prepared in the lab.
@@
Penning traps employ static confining fields and the non-thermal motion of ions in the trap is a rigid body rotation that is not readily shared with the thermal degrees of freedom. Therefore, the implementation of Doppler cooling of rotating ion ensembles in Penning traps is relatively straightforward, although Doppler shifts due the ion crystal rotation can produce temperatures that are slightly elevated from standard Doppler cooling limits [353]. Three-dimensional crystals with $N\sim 5\times 10^{5}$ ions have been Doppler cooled in Penning traps to milliKelvin temperatures ($\Gamma>200$), where a body centered cubic (bcc) lattice structure in the crystal’s interior was observed [343, 168]. Doppler cooling has been demonstrated to be effective for larger ensembles: Nonneutral plasmas of $N\sim 10^{8}$ Mg+ ions in Penning-Malmberg traps have been prepared at temperatures less than 50 mK [4, 3]. Here, the ions are strongly correlated ($\Gamma\sim 20$) but above the temperature for crystal formation. The lack of crystal formation is not understood. Its origin is probably due to the scaling with $N$ of heating processes. For example, the potential energy stored in the ion crystal increases with $N$ and in Penning traps small imperfections in the construction of the trap can drive a slow radial expansion of the ion crystal that converts potential energy to kinetic energy [105, 172].
@@
Cooling large ion ensembles in rf traps is challenging due to the time-dependent potential. In fact, the rf-driven micromoton can be converted into thermal motion by the Coulomb interactions, heating the crystal. This heating rate increases as the temperature is lowered and the number of ions increases, giving rise to chaotic dynamics and limiting the cooling efficiency [29, 286, 258, 280]. Therefore, the formation of large ion crystals in rf traps requires large laser cooling powers. Despite these difficulties, large crystals composed of $10^{4}$ up to $10^{5}$ ions have been prepared in rf traps [90, 158, 253]. It has been observed that once the ion crystal forms, the rf heating can be significantly reduced. A lower limit on the Coulomb coupling parameter of $\Gamma\gtrsim 70$ has been estimated from the observed structure of these large ion crystals in rf traps. Direct measurements of the ion temperature are challenging with 3D crystals in rf traps because of the large driven micromotion.